# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ C.5 Improving the Efficiency of TFB
ACC ($\uparrow$) MLE - 68.99$\pm$0.58 69.10$\pm$2.84 85.65$\pm$0.92 74.53$\pm$0.66 81.52$\pm$0.25 86.53$\pm$0.28
 + TFB (Ours) 10 69.83$\pm$1.02 68.13$\pm$1.03 86.21$\pm$0.90 75.95$\pm$0.34 82.80$\pm$0.35 87.66$\pm$0.35
 + LL TFB 10 70.94$\pm$0.33 68.88$\pm$2.24 85.16$\pm$0.57 75.14$\pm$1.38 83.18$\pm$0.46 87.01$\pm$0.09
 + LL TFB 100 70.83$\pm$0.19 69.14$\pm$2.61 85.48$\pm$0.60 75.22$\pm$0.69 83.41$\pm$0.38 87.56$\pm$0.09
MAP - 68.62$\pm$0.71 67.59$\pm$0.40 86.55$\pm$0.55 75.61$\pm$0.71 81.38$\pm$0.65 86.50$\pm$0.41
 + TFB (Ours) 10 69.17$\pm$1.08 67.68$\pm$1.73 85.86$\pm$0.37 75.87$\pm$0.40 83.07$\pm$0.61 87.74$\pm$0.23
 + LL TFB 10 71.19$\pm$0.54 68.75$\pm$2.23 85.09$\pm$0.64 74.73$\pm$1.13 83.04$\pm$0.48 87.21$\pm$0.01
 + LL TFB 100 70.75$\pm$0.14 69.14$\pm$2.61 85.42$\pm$0.66 75.19$\pm$0.57 83.41$\pm$0.38 87.50$\pm$0.20
BLoB 10 69.07$\pm$0.34 68.81$\pm$1.09 85.56$\pm$0.35 73.69$\pm$0.17 81.52$\pm$0.74 86.99$\pm$0.24
BLoB-Mean - 72.15$\pm$0.17 69.56$\pm$0.91 86.31$\pm$0.37 75.47$\pm$1.36 82.53$\pm$0.74 86.69$\pm$0.08
 + TFB (Ours) 10 69.94$\pm$1.68 70.72$\pm$2.25 86.74$\pm$0.97 73.13$\pm$2.38 83.13$\pm$0.76 86.36$\pm$0.26
 + LL TFB 10 68.48$\pm$1.59 71.61$\pm$1.57 86.00$\pm$0.33 75.00$\pm$0.58 83.63$\pm$0.46 86.49$\pm$0.53
 + LL TFB 100 70.45$\pm$1.01 72.92$\pm$2.56 86.98$\pm$0.91 75.90$\pm$0.33 84.15$\pm$0.32 87.45$\pm$0.48
ECE ($\downarrow$) MLE - 29.83$\pm$0.58 29.00$\pm$1.97 13.12$\pm$1.39 20.62$\pm$0.74 12.55$\pm$0.46 3.18$\pm$0.09
 + TFB (Ours) 10 16.26$\pm$0.36 6.93$\pm$1.43 5.82$\pm$0.87 8.78$\pm$0.84 4.60$\pm$0.62 2.30$\pm$0.50
 + LL TFB 10 20.85$\pm$1.98 14.14$\pm$4.77 4.15$\pm$0.75 11.71$\pm$1.42 4.28$\pm$1.24 1.88$\pm$0.13
 + LL TFB 100 20.66$\pm$2.05 13.81$\pm$4.85 5.54$\pm$1.36 11.03$\pm$1.41 4.25$\pm$0.83 1.49$\pm$0.10
MAP - 29.76$\pm$0.87 29.42$\pm$0.68 12.07$\pm$0.55 23.07$\pm$0.14 13.26$\pm$0.82 3.16$\pm$0.23
 + TFB (Ours) 10 11.72$\pm$0.56 6.07$\pm$1.89 6.99$\pm$0.96 5.21$\pm$0.86 3.82$\pm$0.60 2.65$\pm$0.30
 + LL TFB 10 17.03$\pm$2.51 13.18$\pm$4.16 4.51$\pm$0.38 8.32$\pm$1.26 4.31$\pm$1.58 2.36$\pm$0.21
 + LL TFB 100 16.34$\pm$2.39 12.60$\pm$4.65 6.53$\pm$1.55 6.49$\pm$1.57 4.56$\pm$0.65 2.01$\pm$0.18
BLoB 10 9.35$\pm$1.37 9.59$\pm$1.88 3.64$\pm$0.53 3.01$\pm$0.12 3.77$\pm$1.47 1.41$\pm$0.19
BLoB-Mean - 18.63$\pm$0.31 22.51$\pm$0.93 9.64$\pm$0.60 11.58$\pm$1.24 8.65$\pm$0.98 2.88$\pm$0.07
 + TFB (Ours) 10 6.33$\pm$1.04 5.77$\pm$0.32 3.03$\pm$0.43 4.07$\pm$1.65 5.94$\pm$0.46 5.37$\pm$0.44
 + LL TFB 10 6.36$\pm$1.51 7.35$\pm$1.87 4.18$\pm$0.50 3.31$\pm$0.59 6.30$\pm$0.68 3.91$\pm$0.58
 + LL TFB 100 4.12$\pm$1.36 7.35$\pm$2.22 3.77$\pm$0.41 3.17$\pm$0.27 5.60$\pm$0.72 5.69$\pm$0.53
NLL ($\downarrow$) MLE - 3.17$\pm$0.37 2.85$\pm$0.27 1.17$\pm$0.13 0.95$\pm$0.07 0.73$\pm$0.03 0.32$\pm$0.00
 + TFB (Ours) 10 0.86$\pm$0.06 0.98$\pm$0.02 0.48$\pm$0.04 0.59$\pm$0.01 0.54$\pm$0.02 0.30$\pm$0.00
 + LL TFB 10 1.14$\pm$0.18 1.44$\pm$0.25 0.59$\pm$0.02 0.65$\pm$0.04 0.57$\pm$0.03 0.31$\pm$0.00
 + LL TFB 100 0.85$\pm$0.10 0.98$\pm$0.15 0.48$\pm$0.03 0.60$\pm$0.03 0.50$\pm$0.01 0.30$\pm$0.00
MAP - 2.46$\pm$0.34 2.66$\pm$0.11 0.90$\pm$0.05 1.62$\pm$0.29 0.75$\pm$0.01 0.33$\pm$0.00
 + TFB (Ours) 10 0.72$\pm$0.03 0.96$\pm$0.03 0.50$\pm$0.04 0.55$\pm$0.01 0.53$\pm$0.02 0.30$\pm$0.00
 + LL TFB 10 0.92$\pm$0.12 1.42$\pm$0.24 0.60$\pm$0.02 0.61$\pm$0.03 0.57$\pm$0.03 0.31$\pm$0.00
 + LL TFB 100 0.71$\pm$0.06 0.96$\pm$0.13 0.48$\pm$0.02 0.56$\pm$0.02 0.50$\pm$0.01 0.30$\pm$0.00
BLoB 10 0.63$\pm$0.01 0.78$\pm$0.02 0.40$\pm$0.01 0.54$\pm$0.00 0.50$\pm$0.01 0.31$\pm$0.00
BLoB-Mean - 0.79$\pm$0.01 1.27$\pm$0.02 0.57$\pm$0.03 0.60$\pm$0.03 0.56$\pm$0.00 0.32$\pm$0.01
 + TFB (Ours) 10 0.62$\pm$0.03 0.86$\pm$0.01 0.42$\pm$0.03 0.56$\pm$0.03 0.50$\pm$0.01 0.34$\pm$0.00
 + LL TFB 10 0.63$\pm$0.01 0.90$\pm$0.06 0.44$\pm$0.01 0.54$\pm$0.01 0.48$\pm$0.01 0.33$\pm$0.01
 + LL TFB 100 0.59$\pm$0.01 0.77$\pm$0.03 0.40$\pm$0.02 0.52$\pm$0.01 0.47$\pm$0.01 0.32$\pm$0.01
We report the detailed performance of last-layer TFB (LL TFB) in Table 12. As indicated in the table, with only $N=10$ samples, last-layer Bayesianization provides less effective uncertainty estimation compared to full-model Bayesianization. However, increasing the number of samples to $N=100$ significantly enhances the posterior estimation, allowing last-layer Bayesianization to achieve better accuracy. This improvement further allows it to outperform full-model Bayesianization in terms of NLL across most datasets.

### v2

@@ F.10 Additional Results on Llama2-7B
ACC ($\uparrow$) MLE - 77.87$\pm$0.54 81.08$\pm$0.48 91.67$\pm$0.36 82.30$\pm$0.53 87.90$\pm$0.87 89.58$\pm$0.26
 + TFB 10 77.44$\pm$0.30 82.53$\pm$1.00 91.33$\pm$0.37 82.53$\pm$0.56 88.53$\pm$0.57 89.75$\pm$0.25
 + LL TFB 10 76.96$\pm$0.46 82.00$\pm$0.40 90.97$\pm$0.34 82.67$\pm$0.49 87.80$\pm$1.07 89.62$\pm$0.12
 + LL TFB 100 77.39$\pm$0.32 82.13$\pm$0.82 91.33$\pm$0.37 82.61$\pm$0.55 87.80$\pm$0.91 89.66$\pm$0.28
MAP - 76.90$\pm$0.97 81.08$\pm$2.48 91.61$\pm$0.44 82.59$\pm$0.28 85.73$\pm$0.19 90.09$\pm$0.28
 + TFB (Ours) 10 76.43$\pm$0.72 82.80$\pm$1.42 91.39$\pm$0.37 82.64$\pm$0.58 86.00$\pm$0.16 89.96$\pm$0.18
 + LL TFB 10 76.35$\pm$0.89 83.07$\pm$1.97 91.15$\pm$0.52 82.27$\pm$0.53 85.27$\pm$0.19 90.09$\pm$0.20
 + LL TFB 100 76.72$\pm$0.77 83.07$\pm$2.12 91.15$\pm$0.60 82.53$\pm$0.33 85.60$\pm$0.16 90.02$\pm$0.14
BLoB 10 76.45$\pm$0.37 82.32$\pm$1.15 91.14$\pm$0.54 82.01$\pm$0.56 87.57$\pm$0.21 89.65$\pm$0.15
BLoB-Mean - 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05
 + TFB (Ours) 10 77.81$\pm$0.36 83.33$\pm$0.19 91.76$\pm$0.48 83.81$\pm$0.39 87.80$\pm$0.16 90.11$\pm$0.28
 + LL TFB 10 77.57$\pm$1.02 82.80$\pm$0.33 91.45$\pm$0.54 83.23$\pm$0.57 88.33$\pm$0.09 89.85$\pm$0.13
 + LL TFB 100 77.60$\pm$0.62 83.33$\pm$0.82 91.39$\pm$0.60 83.63$\pm$0.62 87.60$\pm$0.43 90.03$\pm$0.03
ECE ($\downarrow$) MLE - 17.02$\pm$0.46 16.35$\pm$0.68 7.00$\pm$0.53 13.83$\pm$0.65 9.77$\pm$0.81 8.69$\pm$0.21
 + TFB (Ours) 10 12.98$\pm$0.37 11.63$\pm$0.68 5.14$\pm$0.14 10.01$\pm$0.70 7.20$\pm$0.47 7.39$\pm$0.26
 + LL TFB 10 14.42$\pm$0.41 13.86$\pm$0.45 6.92$\pm$0.62 10.32$\pm$0.90 8.56$\pm$0.96 7.52$\pm$0.12
 + LL TFB 100 13.45$\pm$0.30 13.17$\pm$0.62 6.84$\pm$0.67 10.76$\pm$0.88 8.68$\pm$0.60 7.46$\pm$0.10
MAP - 18.71$\pm$0.74 15.77$\pm$1.60 6.62$\pm$0.64 14.26$\pm$0.92 12.19$\pm$0.55 8.40$\pm$0.25
 + TFB (Ours) 10 14.95$\pm$0.65 11.27$\pm$2.53 5.76$\pm$0.63 10.97$\pm$1.19 9.70$\pm$0.69 6.86$\pm$0.31
 + LL TFB 10 16.03$\pm$0.64 12.72$\pm$1.33 6.54$\pm$0.68 12.06$\pm$1.09 11.36$\pm$0.34 7.51$\pm$0.23
 + LL TFB 100 15.56$\pm$0.97 12.84$\pm$2.17 6.38$\pm$0.66 11.80$\pm$1.14 11.22$\pm$0.38 7.30$\pm$0.41
BLoB 10 9.93$\pm$0.22 5.41$\pm$1.17 2.70$\pm$0.87 4.28$\pm$0.64 2.91$\pm$0.92 2.58$\pm$0.25
BLoB-Mean - 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29
 + TFB (Ours) 10 8.16$\pm$0.48 6.48$\pm$0.36 2.44$\pm$0.50 3.83$\pm$0.43 2.67$\pm$0.18 3.10$\pm$0.59
 + LL TFB 10 9.68$\pm$0.70 7.20$\pm$0.91 3.01$\pm$0.66 3.94$\pm$0.78 3.33$\pm$0.93 2.96$\pm$0.30
 + LL TFB 100 8.88$\pm$0.32 6.47$\pm$1.55 2.84$\pm$0.50 3.40$\pm$0.82 3.70$\pm$0.27 2.51$\pm$0.46
NLL ($\downarrow$) MLE - 0.88$\pm$0.04 1.20$\pm$0.11 0.46$\pm$0.04 0.68$\pm$0.01 0.61$\pm$0.06 0.52$\pm$0.01
 + TFB (Ours) 10 0.68$\pm$0.03 0.85$\pm$0.02 0.33$\pm$0.03 0.53$\pm$0.01 0.46$\pm$0.04 0.42$\pm$0.00
 + LL TFB 10 0.70$\pm$0.02 0.96$\pm$0.12 0.41$\pm$0.06 0.53$\pm$0.02 0.50$\pm$0.06 0.42$\pm$0.01
 + LL TFB 100 0.66$\pm$0.02 0.84$\pm$0.08 0.39$\pm$0.07 0.53$\pm$0.02 0.49$\pm$0.05 0.40$\pm$0.00
MAP - 0.99$\pm$0.07 1.12$\pm$0.23 0.46$\pm$0.03 0.74$\pm$0.07 0.79$\pm$0.02 0.52$\pm$0.01
 + TFB (Ours) 10 0.77$\pm$0.05 0.80$\pm$0.15 0.38$\pm$0.03 0.57$\pm$0.05 0.61$\pm$0.03 0.40$\pm$0.01
 + LL TFB 10 0.80$\pm$0.07 0.88$\pm$0.19 0.43$\pm$0.02 0.60$\pm$0.05 0.65$\pm$0.01 0.43$\pm$0.02
 + LL TFB 100 0.77$\pm$0.06 0.86$\pm$0.18 0.41$\pm$0.02 0.57$\pm$0.04 0.63$\pm$0.02 0.40$\pm$0.03
BLoB 10 0.58$\pm$0.00 0.51$\pm$0.03 0.23$\pm$0.01 0.43$\pm$0.01 0.34$\pm$0.01 0.26$\pm$0.01
BLoB-Mean - 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02
 + TFB 10 0.55$\pm$0.01 0.53$\pm$0.04 0.23$\pm$0.02 0.40$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01
 + LL TFB 10 0.56$\pm$0.02 0.60$\pm$0.05 0.26$\pm$0.02 0.41$\pm$0.01 0.33$\pm$0.01 0.27$\pm$0.01
 + LL TFB 100 0.53$\pm$0.01 0.54$\pm$0.04 0.24$\pm$0.01 0.39$\pm$0.01 0.31$\pm$0.01 0.26$\pm$0.01
Table 15: Performance of different methods applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. “TF?” denotes whether a method is Training-Free. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We sample $N=10$ during inference in all sampling-based methods including BLoB [61] and TFB. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. For TFB, the anchor dataset ${\mathcal{D}}$ is set to a randomly sampled subset of the original training set, the performance evaluation metric $l$ is set to accuracy, and the performance drop tolerance is set adaptively to 1% or 0.5% based on whether the given mean overfits. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
[…]
Small Shift Large Shift
WG-S ARC-C ARC-E WG-M OBQA BoolQ ARC-C ARC-E Chem Phy